OpenStudy (anonymous):

Question 3 (Multiple Choice Worth 4 points) [4.06]What is the equation of the line that passes through the points (−1, 3) and (1, 11)? 4x − y = 1 4x + y = 1 x − 4y = −13 4x − y = −7

OpenStudy (anonymous):

@mahmit2012

OpenStudy (anonymous):

@satellite73 @sauravshakya @mathteacher1729 @KerenNicole @dmezzullo @hba

OpenStudy (anonymous):

Question 23 (Multiple Choice Worth 2 points) [3.05] The set of ordered pairs below defines a relation. In each ordered pair (x, y), x represents the daily wages of a plumber, and y represents the daily wages of her assistant. {(144, 72), (171, 84), (198, 84), (171, 96), (144, 72)} What is the difference between the values of the least and greatest elements in the domain of this relation? 24 54 72 126

OpenStudy (hba):

Do You Know How To Find The Slope ?

OpenStudy (anonymous):

this one too plz

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

y2 -y1 / x2-x1

OpenStudy (mathteacher1729):

Is this from a FLVS graded test?

OpenStudy (hba):

Ok So In The First one You Have To Find The Slope and then put it in (y-y1)=m(x-x1)

OpenStudy (anonymous):

the slope is 3/8?

OpenStudy (anonymous):

what is the answer

OpenStudy (anonymous):

plz for both

OpenStudy (anonymous):

@henpen

OpenStudy (anonymous):

@georgejafacules @sauravshakya @satellite73 @sara12345 @stugglechickk

OpenStudy (anonymous):

Find slope. \[Slope =\frac{ y2-y1 }{ x2-x1 } = \frac{ 11-3 }{ 1 - (-1) } = \frac{ 11-3 }{ 1 + 1 } =\frac{ 8 }{2 } = \frac{ 4 }{ 1 }=4\]Slope is 4

OpenStudy (anonymous):

yeah the equation is y - 11=4(x-1)

OpenStudy (anonymous):

I found y=4x-7

OpenStudy (anonymous):

what is the answer for both Please

OpenStudy (anonymous):

But that is not one of the options

OpenStudy (anonymous):

check in the question for the options

OpenStudy (anonymous):

thanx

OpenStudy (anonymous):

it is only 4x-y=-7

OpenStudy (anonymous):

so it is not an option

OpenStudy (anonymous):

@Fall12-13 @sauravshakya help

OpenStudy (anonymous):

i just need the answer

OpenStudy (anonymous):

@TuringTest @chak33056 @ganeshie8 @ParthKohli

OpenStudy (anonymous):

Help